# No Metric to Rule Them All: Toward Principled Evaluations of Graph-Learning Datasets

### v2

@@ 2 The Rings Framework
[…]
@@ 2.2 Performance Separability
A tuned model $\operatorname{\mathcal{M}}$ is a triple $\operatorname{\mathcal{M}}\coloneqq(\operatorname{\mathcal{D}},\operatorname{\mathcal{A}},\operatorname{\theta})$, where $\operatorname{\mathcal{D}}$ represents the dataset and associated task, $\operatorname{\mathcal{A}}$ is the architecture used during training, and $\operatorname{\theta}$ denotes the tuned parameters for the specific architecture.
@@ 2.3 Mode Complementarity
$D_{S}\coloneqq D_{\overline{S}}/\text{diam}(D_{\overline{S}})$  for $\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ \leavevmode\nobreak\ }D_{\overline{S}}\coloneqq\operatorname{\mathcal{L}}_{\operatorname{d}_{S}}(G,X)\;,$ (7)
$D_{F}\coloneqq D_{\overline{F}}/\text{diam}(D_{\overline{F}})$  for $\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ \leavevmode\nobreak\ }D_{\overline{F}}\coloneqq\operatorname{\mathcal{L}}_{\operatorname{d}_{F}}(G,X)\,,$ (8)
@@
[…]
@@ 3 Related Work
AIDS cf $>$ cg $>$ rg $>$ o $>$ eg $>$ rf cf/cg/rg $>$ eg/o $>$ rf uninformative uninformative --
COLLAB o $>$ cg/rg $>$ eg $>$ cf/rf o $>$ cg/rg $>$ eg $>$ rf $>$ cf informative informative ++
DD rg $>$ cg $>$ cf $>$ eg/o $>$ rf rg $>$ cf/cg/eg/o/rf uninformative uninformative --
Enzymes eg $>$ cg/o $>$ rg $>$ cf $>$ rf eg $>$ o $>$ cg $>$ rg $>$ cf $>$ rf uninformative informative -
IMDB-B cf/cg/eg/o/rf/rg o $>$ cg $>$ eg/rg $>$ rf $>$ cf (un)informative (un)informative $\circ$
IMDB-M cg/eg/o/rg $>$ cf $>$ rf o $>$ cg/eg/rg $>$ cf $>$ rf (un)informative informative +
MUTAG cf/cg/o/rg $>$ eg $>$ rf cf/o $>$ cg/eg/rf/rg (un)informative uninformative -
MolHIV o $>$ cf/cg/rg $>$ rf $>$ eg o $>$ cg $>$ cf $>$ rg $>$ eg $>$ rf informative informative ++
NCI1 o $>$ cg $>$ cf $>$ rg $>$ eg $>$ rf o $>$ cg $>$ cf $>$ rg $>$ eg/rf informative informative ++
Peptides o $>$ cg $>$ rg $>$ eg $>$ cf $>$ rf cg $>$ o $>$ rg $>$ eg $>$ cf $>$ rf (un)informative informative +
Proteins cf $>$ cg/rf $>$ eg/o/rg cf/cg/eg/o/rf/rg uninformative uninformative --
Reddit-B cf $>$ rg $>$ o $>$ cg $>$ eg/rf rg $>$ cf $>$ o $>$ cg/eg/rf uninformative uninformative --
Reddit-M rf $>$ rg $>$ o $>$ eg rg $>$ rf $>$ o $>$ eg uninformative uninformative --
Table 1: Measuring performance separability between different versions of the same dataset. To quantify the conclusions from Figure 3 and further account for performance distributions (Definition 2.7), we use permutation tests with 10 0001000010\,00010 000 random permutations and the Kolmogorov-Smirnov (KS) statistic as our test statistic at an $\alpha$-level of $0.01$, Bonferroni-corrected for multiple hypothesis testing within each individual dataset. Here, for sets $S_{1}$ and $S_{2}$, $S_{1}>S_{2}$ denotes that the elements in the set $S_{1}$ separably outperform the elements in the set $S_{2}$ (i.e., the pairwise distances between $s_{1}$ and $s_{2}$ are statistically significantly different for all $s_{1}\in S_{1}$ and $s_{2}\in S_{2}$), and we represent the sets in condensed notation, concatenating elements with “/ ”. We see that original datasets are often separably outperformed by their perturbed variants.
@@ 4.1 Performance Separability (P1)
Figure 4: Comparing levels of mode complementarity across different versions of the same dataset. We show the mean (dot) and $95$th percentile intervals (bars) of complementarity scores for the original version as well as $4$ deterministic perturbations (top) and $4$ randomized perturbations (bottom) of $13$ graph-learning datasets, computed with $t=1$ diffusion steps. Black horizontal lines indicate mean mode complementarities of the original dataset, and the silver horizontal line marks the $0.5$ threshold relevant for assessing mode diversity. Note that $\operatorname{\gamma}_{\text{eg}}=1-\operatorname{\gamma}_{\text{cg}}$ and $\operatorname{\gamma}_{\text{ef}}=1-\operatorname{\gamma}_{\text{cf}}$ by Theorem 2.15.
@@ 4.2 Mode Complementarity (P2)
[…]
@@ 4.3 Dataset Taxonomy
[…]
@@
[…]
@@
Taxonomy, categories, and interpretation. Our taxonomy is built on our measures of performance separability and structural diversity (derived from mode complementarity). Here, we emphasize structural diversity over feature diversity because the structural mode is characteristic of the graph-learning setting, acknowledging that datasets with high feature diversity could still be useful as benchmarks in settings outside of graph learning. Here, we call the value of a measure high if it is judged as $\circ$, $+$, or +++++ + (see Tables 1 and 2), and low otherwise. In the following, we elaborate on the definition and interpretation of each action category.
[…]
@@ A.2 Perturbations
[…]
@@ A.3.1 Comparators
$||D||_{p,q}=\left(\frac{\sum_{i=1}^{n}\left(\sum_{j=1}^{n}\lvert\operatorname{d}(i,j)\rvert^{p}\right)^{q/p}}{n(n-1)}\right)^{1/q}\;.$ (13)
@@ A.3.2 Complementarity under Perturbation
=$\overset{\operatorname{\operatorname{d}_{F}}\leq 1}{=}\frac{1}{n^{2}-n}\left(n^{2}-n-\sum_{i=1}^{n}\sum_{j\neq i}^{n}\operatorname{\operatorname{d}_{F}}(i,j)\right)$
@@
[…]
@@ B.2 Metric Choices
Figure 6: Mode complementarity under different graph distances and varying density. We show the evolution of mode complementarity for $5$ graphs on $10$ nodes with normally distributed one-dimensional features as we randomly add edges to increase the density from the minimum for a connected graph ($0.2$) to the maximum ($1.0$), for different choices of graph distance $\operatorname{d}_{S}$. For the diffusion distance and the heat-kernel distance, which depend on the parameter $t$, we show the complementarity at $t=1$. The diffusion distance exhibits the smoothest behavior.
@@ B.2.1 Structural Metrics
Definition B.2 (Diffusion Distance [$\operatorname{\operatorname{d}_{S_{t}}}$] (Coifman & Lafon, 2006)).
@@
with $+$ denoting the Moore-Penrose pseudoinverse.
@@ B.2.2 Feature Metrics
Definition B.6 (Euclidean Distance [$\operatorname{d}_{F}$]).
@@ Appendix C Extended Dataset Descriptions
$N$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
@@
Table 3: Basic statistics of our $13$ evaluated graph-learning datasets. We show the number of graphs ($N$) as well as the mean and standard deviation of the number of nodes ($n$), the number of edges ($m$), the degree ($\delta$), the density ($\rho$), and mode complementarity at $t\in\{1,10\}$ ($\gamma_{1}$, $\gamma_{10}$).
@@ C.2.2 Larger Chemical Structures
DD, ENZYMES, and Proteins (Proteins-full) are all protein datasets. Each dataset draws upon the framework introduced by (Borgwardt et al., 2005) to model 3D structures of folded proteins as graphs, with nodes representing amino acids and edges connecting amino acids within a given proximity measured in Angstroms. Both the Proteins-full database introduced by (Borgwardt et al., 2005) and the DD database used by (Shervashidze et al., 2011) are based on data from Dobson & Doig (2003), which classifies 1 17811781\,1781 178 proteins as either enzymes ($59\%$) or non-enzymes ($41\%$). Thus, these two datasets are binary-classification datasets.
Proteins-full only uses 1 12811281\,1281 128 of the 1 17811781\,1781 178 proteins from Dobson & Doig (2003) but encodes $29$ descriptive node features, whereas DD encodes none (Kersting et al., 2016; Morris et al., 2020). This distinction is responsible for Proteins-full’s drastically higher feature diversity score and standard deviation, as seen in Table 2.
@@ D.2.2 Performance-Separability Results
For the judgment depicted in the Evaluation column of Table 1, we score datasets on a scale from 00 to $5$ as $1.5\cdot S+1\cdot F$, where $S,F\in\{0,1,2\}$, mapping “uninformative” to 00, “(un)informative” to $1$, and informative to $2$, and giving the structural mode higher weight than the feature mode because we are dealing with graph-learning datasets. The final evaluation results from binning the scores as
[…]
@@ D.2.3 Performance-Separability Evaluation
AIDS cf $>$ cg $>$ rg $>$ o $>$ eg $>$ rf cf/cg/eg/o/rg $>$ rf uninformative uninformative --
COLLAB o $>$ cg/rg $>$ eg $>$ cf/rf o $>$ cg/rg $>$ eg $>$ rf $>$ cf informative informative ++
DD rg $>$ cg $>$ cf $>$ eg/o $>$ rf rg $>$ cf/cg/eg/o/rf uninformative uninformative --
Enzymes eg $>$ cg/o $>$ rg $>$ cf $>$ rf eg/o $>$ cg $>$ rg $>$ cf $>$ rf uninformative informative -
IMDB-B cf/cg/eg/o/rf/rg cg/o $>$ eg/rg $>$ cf/rf uninformative (un)informative --
IMDB-M cg/eg/o/rg $>$ cf $>$ rf o $>$ eg/rg $>$ cg $>$ cf $>$ rf (un)informative informative +
MUTAG cf/cg/eg/o/rg $>$ rf cf/cg/eg/o/rf/rg uninformative uninformative --
MolHIV o $>$ cf/cg/rg $>$ rf $>$ eg o $>$ cf/cg $>$ rg $>$ eg $>$ rf informative informative ++
NCI1 o $>$ cg $>$ cf $>$ rg $>$ eg/rf o $>$ cg $>$ cf $>$ rg $>$ eg/rf informative informative ++
Peptides o $>$ cg $>$ rg $>$ eg $>$ cf $>$ rf cg $>$ o $>$ rg $>$ eg $>$ cf $>$ rf (un)informative informative +
Proteins cf $>$ cg/eg/o/rf/rg cf/cg/eg/o/rf/rg uninformative uninformative --
Reddit-B cf $>$ rg $>$ o $>$ cg $>$ eg/rf rg $>$ cf $>$ o $>$ cg/eg/rf uninformative uninformative --
Reddit-M rf/rg $>$ o $>$ eg rg $>$ rf $>$ o $>$ eg uninformative uninformative --
Table 10: Measuring performance separability between different versions of the same dataset: Wilcoxon rank-sum test. Supplementing Table 1, we use the Wilcoxon rank-sum as a test statistic to replace the KS statistic in our permutation tests, with an otherwise identical setup (10 0001000010\,00010 000 random permutations; $\alpha$-level of $0.01$; Bonferroni correction within each dataset). The results are substantively the same, with the exceptions that IMDB-B gets downgraded from fair to poor, and MUTAG gets downgraded from poor to very poor.
@@
AIDS cf $>$ cg $>$ rg $>$ eg/o $>$ rf cf/cg/rg $>$ eg/o $>$ rf uninformative uninformative --
COLLAB o $>$ cg/rg $>$ eg $>$ cf/rf o $>$ cg/rg $>$ eg $>$ rf $>$ cf informative informative ++
DD rg $>$ cg $>$ cf $>$ eg/o $>$ rf rg $>$ cf/cg/eg/o/rf uninformative uninformative --
Enzymes eg $>$ cg/o $>$ rg $>$ cf $>$ rf eg/o $>$ cg $>$ rg $>$ cf $>$ rf uninformative informative -
IMDB-B cf/cg/eg/o/rf/rg o $>$ cg $>$ eg/rg $>$ cf/rf (un)informative (un)informative $\circ$
IMDB-M cg/eg/o/rg $>$ cf $>$ rf o $>$ cg/eg/rg $>$ cf $>$ rf (un)informative informative +
MUTAG cf/cg/o/rg $>$ eg $>$ rf cf/o $>$ cg/eg/rf/rg (un)informative uninformative -
MolHIV o $>$ cf/cg/rg $>$ rf $>$ eg o $>$ cf/cg $>$ rg $>$ eg $>$ rf informative informative ++
NCI1 o $>$ cg $>$ cf $>$ rg $>$ eg $>$ rf o $>$ cg $>$ cf $>$ rg $>$ eg/rf informative informative ++
Peptides o $>$ cg $>$ rg $>$ eg $>$ cf $>$ rf cg $>$ o $>$ rg $>$ eg $>$ cf $>$ rf (un)informative informative +
Proteins cf $>$ cg/eg/o/rf/rg cf/cg/eg/o/rf/rg uninformative uninformative --
Reddit-B cf $>$ o/rg $>$ cg $>$ eg/rf rg $>$ cf $>$ o $>$ cg/eg/rf uninformative uninformative --
Reddit-M rf/rg $>$ o $>$ eg rg $>$ rf $>$ o $>$ eg uninformative uninformative --
Table 11: Measuring performance separability between different versions of the same dataset: $\mathbf{\alpha}\mathbf{\leq}\mathbf{0.005}$. Supplementing Table 1, we test at an $\alpha$-level of $0.005$ instead of $0.01$, using an otherwise identical setup (10 0001000010\,00010 000 random permutations; KS statistic; Bonferroni correction within each dataset). The results are substantively identical to Table 1.

### v3

@@ 2 The Rings Framework
Preliminaries. We work with attributed graphs $(G,X)$, where $G=(V,E)$ is a graph with $n=|V|$ nodes and $m=|E|$ edges, $X\subset\mathds{R}^{k}$ is the space of $k$-dimensional node features, and we assume w.l.o.g. that |X|=n{|X|=n}. For graph-level tasks, we have datasets $\operatorname{\mathcal{D}}=\{G_{1},\dots,G_{N}\}$}, where $N$ is the total number of graphs. Given a set $S$, $2^{S}$ is its power set, and $\binom{S}{\ell}$ is the set of all $\ell$-element subsets of $S$. Multisets are denoted by $\{\{\cdot\}\}$, and the set of positive integers no greater than $\ell$ is written as $\ell]=\{i\in\mathbb{Z}\mid 0<i\leq\ell\}$}.
@@ 2.2 Performance Separability
[…]
@@ 2.3 Mode Complementarity
$D_{S}\coloneqq D_{\overline{S}}/\text{diam}(D_{\overline{S}})$  $\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ \leavevmode\nobreak\ }D_{\overline{S}}\coloneqq\operatorname{\mathcal{L}}_{\operatorname{d}_{S}}(G,X)\;,$ (7)
$D_{F}\coloneqq D_{\overline{F}}/\text{diam}(D_{\overline{F}})$  $\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ \leavevmode\nobreak\ }D_{\overline{F}}\coloneqq\operatorname{\mathcal{L}}_{\operatorname{d}_{F}}(G,X)\,$, (8)
@@
Fix an attributed graph $(G,X)$ and corresponding distances $\operatorname{d}_{S}$$\operatorname{d}_{F}$ for lifting each mode into a metric space. For $\ast\in\{\text{f},\text{g}\}$, Definition 2.12 of $\operatorname{\gamma}$ yields the equivalence
@@ 3 Related Work
AIDS cf >> cg >> rg >> o >> eg >> rf cf/cg/rg >> eg/o >> rf uninformative uninformative --
COLLAB o >> cg/rg >> eg >> cf/rf o >> cg/rg >> eg >> rf >> cf informative informative ++
$D$ rg >> cg >> cf >> eg/o >> rf rg >> cf/cg/eg/o/rf uninformative uninformative --
Enzymes eg >> cg/o >> rg >> cf >> rf eg >> o >> cg >> rg >> cf >> rf uninformative informative -
IMDB-B cf/cg/eg/o/rf/rg o >> cg >> eg/rg >> rf >> cf (un)informative (un)informative $\circ$
IMDB-M cg/eg/o/rg >> cf >> rf o >> cg/eg/rg >> cf >> rf (un)informative informative +
MUTAG cf/cg/o/rg >> eg >> rf cf/o >> cg/eg/rf/rg (un)informative uninformative -
MolHIV o >> cf/cg/rg >> rf >> eg o >> cg >> cf >> rg >> eg >> rf informative informative ++
NCI1 o >> cg >> cf >> rg >> eg >> rf o >> cg >> cf >> rg >> eg/rf informative informative ++
Peptides o >> cg >> rg >> eg >> cf >> rf cg >> o >> rg >> eg >> cf >> rf (un)informative informative +
Proteins cf >> cg/rf >> eg/o/rg cf/cg/eg/o/rf/rg uninformative uninformative --
Reddit-B cf >> rg >> o >> cg >> eg/rf rg >> cf >> o >> cg/eg/rf uninformative uninformative --
Reddit-M rf >> rg >> o >> eg rg >> rf >> o >> eg uninformative uninformative --
Table 1: Measuring performance separability between different versions of the same dataset. To quantify the conclusions from Figure 3 and further account for performance distributions (Definition 2.7), we use permutation tests with 10 00010\,000 random permutations and the Kolmogorov-Smirnov (KS) statistic as our test statistic at an $\alpha$-level of $0.01$, Bonferroni-corrected for multiple hypothesis testing within each individual dataset. Here, for sets $S_{1}$ and $S_{2}$, $S_{1}>S_{2}$ denotes that the elements in the set $S_{1}$ separably outperform the elements in the set $S_{2}$ (i.e., the pairwise distances between $s_{1}$ and $s_{2}$ are statistically significantly different for all $s_{1}\in S_{1}$ and $s_{2}\in S_{2}$), and we represent the sets in condensed notation, concatenating elements with “/ ”. We see that original datasets are often separably outperformed by their perturbed variants.
@@ 4.1 Performance Separability (P1)
Figure 4: Comparing levels of mode complementarity across different versions of the same dataset. We show the mean (dot) and $95$th percentile intervals (bars) of complementarity scores for the original version as well as 44 deterministic perturbations (top) and $4$ randomized perturbations (bottom) of $13$ graph-learning datasets, computed with $t=1$ diffusion steps. Black horizontal lines indicate mean mode complementarities of the original dataset, and the silver horizontal line marks the $0.5$ threshold relevant for assessing mode diversity. Note that $\operatorname{\gamma}_{\text{eg}}=1-\operatorname{\gamma}_{\text{cg}}$ and $\operatorname{\gamma}_{\text{ef}}=1-\operatorname{\gamma}_{\text{cf}}$ by Theorem 2.15.
@@ 4.2 Mode Complementarity (P2)
Measurement. We evaluate mode complementarity for the original as well as $4$ fixed and $4$ randomized mode perturbations: {complete, empty, random, shuffled} $\times$ {graph, features} (see Appendix D for detailed descriptions). To measure mode complementarity (and the mode diversity derived from it), for each graph in a given dataset perturbation, we instantiate Definition 2.12 using the Euclidean distance as our feature distance, the diffusion distance (see Appendix B) for a number of diffusion steps $t\in[10]$ as our graph distance, and the $L_{1,1}$ norm as our comparator (see Appendix B for examples of other choices).
@@ 4.3 Dataset Taxonomy
From our observations on performance separability (P1) and mode complementarity (P2), we can distill the following actionable dataset taxonomy, where we note evidence from performance separability ($\dagger$) and mode complementarity ($\ddagger$}):
@@
Keep ($\dagger|{\ddagger}$) MolHIV, NCI1, Peptides
Realign ($\ddagger$}) AIDS, DD, MUTAG, Reddit-B, Reddit-M
Deprecate ($\ddagger$}) COLLAB, IMDB-B, IMDB-M
Deprecate ($\dagger|{\ddagger}$) Enzymes, Proteins
@@
Taxonomy, categories, and interpretation. Our taxonomy is built on our measures of performance separability and structural diversity (derived from mode complementarity). Here, we emphasize structural diversity over feature diversity because the structural mode is characteristic of the graph-learning setting, acknowledging that datasets with high feature diversity could still be useful as benchmarks in settings outside of graph learning. Here, we call the value of a measure high if it is judged as $\circ$, ++, or ++++ (see Tables 1 and 2), and low otherwise. In the following, we elaborate on the definition and interpretation of each action category.
Keep ($\dagger|{\ddagger}$). This category contains datasets with high performance separability and high structural diversity. These datasets constitute valuable graph-learning benchmarks.
Realign ($\ddagger$}). We call a dataset misaligned if it exhibits low performance separability but high structural diversity. Realignment, then, collectively denotes several potential operations, including changing the prediction targets (e.g., using different categories to classify discussion threads in Reddit-B or Reddit-M) or changing the prediction task (e.g., moving from a graph-level to a node-level task).
Deprecate ($\ddagger$}). This category contains datasets with high performance separability but low structural diversity. Although both structure and features may be needed to achieve state-of-the-art performance, these datasets do not contain interesting structural variation, and as such, they do little to probe the capabilities of graph-learning models.
Deprecate ($\dagger|{\ddagger}$). This category contains datasets that exhibit low structural diversity and low performance separability. These datasets do not currently require the capabilities of graph-learning models to integrate information from structure and features, and they also do not hold sufficient structural diversity to change this via realignment.
@@ A.2 Perturbations
$\operatorname{\mathcal{R}_{F}}_{,1}(X,k)$ $\mapsto\{\mathbf{x_{i}}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{k})\}$ $[\text{random features}]$
$\operatorname{\mathcal{R}_{F}}_{,2}(X)$ $\coloneqq\pi(X)$ $[\text{shuffled features}]$
$\operatorname{\mathcal{R}_{S}}_{,1}(G,p)$ $\coloneqq ER(n,p)$ $[\text{random graph}]$
$\operatorname{\mathcal{R}_{S}}_{,2}(G)$ $\coloneqq(V,\pi(E))$ $[\text{shuffled graph}]$
@@ A.3.1 Comparators
$D||_{p,q}=\left(\frac{\sum_{i=1}^{n}\left(\sum_{j=1}^{n}\lvert\operatorname{d}(i,j)\rvert^{p}\right)^{q/p}}{n(n-1)}\right)^{1/q}\;.$ (13)
@@ A.3.2 Complementarity under Perturbation
$\overset{\operatorname{\operatorname{d}_{F}}\leq 1}{=}\frac{1}{n^{2}-n}\left(n^{2}-n-\sum_{i=1}^{n}\sum_{j\neq i}^{n}\operatorname{\operatorname{d}_{F}}(i,j)\right)$
@@
Let $\operatorname{\operatorname{d}_{S}}$ and $\operatorname{\operatorname{d}_{F}}$ be metrics over the modes of an attributed graph $(G,X)$. We define $(\operatorname{\operatorname{\varphi}_{eg}},\operatorname{\operatorname{d}_{F}})$ and $(\operatorname{\operatorname{\varphi}_{ef}},\operatorname{\operatorname{d}_{S}})$ as dual pairs that can be used to analyze the self-complementarity of a single mode. By construction, for $\ast\in\{\text{f},\text{g}\}$, we have:
@@ B.2 Metric Choices
Figure 6: Mode complementarity under different graph distances and varying density. We show the evolution of mode complementarity for 55 graphs on $10$ nodes with normally distributed one-dimensional features as we randomly add edges to increase the density from the minimum for a connected graph ($0.2$) to the maximum ($1.0$), for different choices of graph distance $\operatorname{d}_{S}$. For the diffusion distance and the heat-kernel distance, which depend on the parameter $t$, we show the complementarity at $t=1$. The diffusion distance exhibits the smoothest behavior.
@@ B.2.1 Structural Metrics
Definition B.2 (Diffusion Distance $\operatorname{\operatorname{d}_{S_{t}}}$] (Coifman & Lafon, 2006)).
@@
with ++ denoting the Moore-Penrose pseudoinverse.
@@ B.2.2 Feature Metrics
Definition B.6 (Euclidean Distance $\operatorname{d}_{F}$]).
@@ Appendix C Extended Dataset Descriptions
NN $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
@@
Table 3: Basic statistics of our $13$ evaluated graph-learning datasets. We show the number of graphs (NN) as well as the mean and standard deviation of the number of nodes ($n$), the number of edges ($m$), the degree ($\delta$), the density ($\rho$), and mode complementarity at $t\in\{1,10\}$ ($\gamma_{1}$, $\gamma_{10}$).
@@ C.2.2 Larger Chemical Structures
DD, ENZYMES, and Proteins (Proteins-full) are all protein datasets. Each dataset draws upon the framework introduced by (Borgwardt et al., 2005) to model 3D structures of folded proteins as graphs, with nodes representing amino acids and edges connecting amino acids within a given proximity measured in Angstroms. Both the Proteins-full database introduced by (Borgwardt et al., 2005) and the DD database used by (Shervashidze et al., 2011) are based on data from Dobson & Doig (2003), which classifies 1 1781\,178 proteins as either enzymes ($59\%$) or non-enzymes ($41\%$). Thus, these two datasets are binary-classification datasets.
Proteins-full only uses 1 1281\,128 of the 1 1781\,178 proteins from Dobson & Doig (2003) but encodes $29$ descriptive node features, whereas DD encodes none (Kersting et al., 2016; Morris et al., 2020). This distinction is responsible for Proteins-full’s drastically higher feature diversity score and standard deviation, as seen in Table 2.
@@ D.2.2 Performance-Separability Results
For the judgment depicted in the Evaluation column of Table 1, we score datasets on a scale from 0 to $5$ as $1.5\cdot S+1\cdot F$, where $S,F\in\{0,1,2\}$, mapping “uninformative” to 0, “(un)informative” to $1$, and informative to $2$, and giving the structural mode higher weight than the feature mode because we are dealing with graph-learning datasets. The final evaluation results from binning the scores as
$[0,1]\mapsto\leavevmode\nobreak\ \texttt{--},\leavevmode\nobreak\ (1,2]\mapsto\leavevmode\nobreak\ \texttt{-},$
$(2,3]\mapsto\leavevmode\nobreak\ \circ,\leavevmode\nobreak\ (3,4]\mapsto\leavevmode\nobreak\ \texttt{+},\leavevmode\nobreak\ (4,5]\mapsto\leavevmode\nobreak\ \texttt{++}.$
@@ D.2.3 Performance-Separability Evaluation
AIDS cf >> cg >> rg >> o >> eg >> rf cf/cg/eg/o/rg >> rf uninformative uninformative --
COLLAB o >> cg/rg >> eg >> cf/rf o >> cg/rg >> eg >> rf >> cf informative informative ++
$D$ rg >> cg >> cf >> eg/o >> rf rg >> cf/cg/eg/o/rf uninformative uninformative --
Enzymes eg >> cg/o >> rg >> cf >> rf eg/o >> cg >> rg >> cf >> rf uninformative informative -
IMDB-B cf/cg/eg/o/rf/rg cg/o >> eg/rg >> cf/rf uninformative (un)informative --
IMDB-M cg/eg/o/rg >> cf >> rf o >> eg/rg >> cg >> cf >> rf (un)informative informative +
MUTAG cf/cg/eg/o/rg >> rf cf/cg/eg/o/rf/rg uninformative uninformative --
MolHIV o >> cf/cg/rg >> rf >> eg o >> cf/cg >> rg >> eg >> rf informative informative ++
NCI1 o >> cg >> cf >> rg >> eg/rf o >> cg >> cf >> rg >> eg/rf informative informative ++
Peptides o >> cg >> rg >> eg >> cf >> rf cg >> o >> rg >> eg >> cf >> rf (un)informative informative +
Proteins cf >> cg/eg/o/rf/rg cf/cg/eg/o/rf/rg uninformative uninformative --
Reddit-B cf >> rg >> o >> cg >> eg/rf rg >> cf >> o >> cg/eg/rf uninformative uninformative --
Reddit-M rf/rg >> o >> eg rg >> rf >> o >> eg uninformative uninformative --
Table 10: Measuring performance separability between different versions of the same dataset: Wilcoxon rank-sum test. Supplementing Table 1, we use the Wilcoxon rank-sum as a test statistic to replace the KS statistic in our permutation tests, with an otherwise identical setup (10 00010\,000 random permutations; $\alpha$-level of $0.01$; Bonferroni correction within each dataset). The results are substantively the same, with the exceptions that IMDB-B gets downgraded from fair to poor, and MUTAG gets downgraded from poor to very poor.
@@
AIDS cf >> cg >> rg >> eg/o >> rf cf/cg/rg >> eg/o >> rf uninformative uninformative --
COLLAB o >> cg/rg >> eg >> cf/rf o >> cg/rg >> eg >> rf >> cf informative informative ++
$D$ rg >> cg >> cf >> eg/o >> rf rg >> cf/cg/eg/o/rf uninformative uninformative --
Enzymes eg >> cg/o >> rg >> cf >> rf eg/o >> cg >> rg >> cf >> rf uninformative informative -
IMDB-B cf/cg/eg/o/rf/rg o >> cg >> eg/rg >> cf/rf (un)informative (un)informative $\circ$
IMDB-M cg/eg/o/rg >> cf >> rf o >> cg/eg/rg >> cf >> rf (un)informative informative +
MUTAG cf/cg/o/rg >> eg >> rf cf/o >> cg/eg/rf/rg (un)informative uninformative -
MolHIV o >> cf/cg/rg >> rf >> eg o >> cf/cg >> rg >> eg >> rf informative informative ++
NCI1 o >> cg >> cf >> rg >> eg >> rf o >> cg >> cf >> rg >> eg/rf informative informative ++
Peptides o >> cg >> rg >> eg >> cf >> rf cg >> o >> rg >> eg >> cf >> rf (un)informative informative +
Proteins cf >> cg/eg/o/rf/rg cf/cg/eg/o/rf/rg uninformative uninformative --
Reddit-B cf >> o/rg >> cg >> eg/rf rg >> cf >> o >> cg/eg/rf uninformative uninformative --
Reddit-M rf/rg >> o >> eg rg >> rf >> o >> eg uninformative uninformative --
Table 11: Measuring performance separability between different versions of the same dataset: $\mathbf{\alpha}\mathbf{\leq}\mathbf{0.005}$. Supplementing Table 1, we test at an $\alpha$-level of $0.005$ instead of $0.01$, using an otherwise identical setup (10 00010\,000 random permutations; KS statistic; Bonferroni correction within each dataset). The results are substantively identical to Table 1.